I have done these problems before but I am stumped.....
The area of a rectangular wall is 240 sq. ft. It's length is 6 feet shorter than three times it's width. Find the length and width of the wall.
I know A=LW and I set it up as w(3w-6)=240
I got answers that don't work. How do I solve this? Have I set it up correctly??
Thanks,
Tammy

Answer by stanbon(75887)

The area of a rectangular wall is 240 sq. ft. It's length is 6 feet shorter than three times it's width. Find the length and width of the wall.
I know A=LW and I set it up as w(3w-6)=240
--------------------------
3w^2-6w = 240
w^2-2w-80 = 0
-------
(w-10)(w+8) = 0
width = 10 ft.
length = 3w-6 = 30-6 = 24 ft
================================
Cheers,
Stan H.
================================